OpenStudy (anonymous):

Solve. - 2/5n = -30 n = Blank

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

\[\frac{ -2 }{ 5 }n = -30\]

OpenStudy (danjs):

multiply both sides by 5 first

OpenStudy (anonymous):

-30 * 5 = 150?

OpenStudy (danjs):

right, that is what the right side becomes, do the same to the left side

OpenStudy (anonymous):

5/1 * -2/5 = 10/5

OpenStudy (danjs):

Remember, you can do whatever operation you want to the equation, you have to just do it on both sides of the equality

OpenStudy (danjs):

\[\frac{ 5 }{ 1 }*\frac{ -2 }{ 5 } = \frac{ -10 }{ 5 } = -2\]

OpenStudy (danjs):

When you multiply, this happens \[\frac{ a }{ b }*\frac{ c }{ d } = \frac{ a*c }{ b*d }\]

OpenStudy (danjs):

that was the purpose of multiplying by 5, to get rid of the fraction

OpenStudy (anonymous):

Ohhh I got one more and Thanks ^_^

OpenStudy (danjs):

-2n = -150 divide by -2 on both sides n = 150/2 = 75

OpenStudy (anonymous):

I'm gonna post the last one

OpenStudy (danjs):

k